OpenStudy (anonymous):

express this amount in dozens. (use deciaml form.) 7.5 cupcakes

OpenStudy (gorv):

1 dozen = 12 things or cupcakes 12 cupcakes= 1dozen 1 cupcakes=1/12 dozen 7.5 cupcakes=(1/12)*7.5=7.5/12=75/120=15/24=5/8=0.625